OpenStudy (cenaida):

Use a system of equations to solve the word problem. The sum of two numbers is 24. Their difference is 14. Find the numbers. A. 5 & 19 B. 7 & 17 C. 10 & 14 D. 6 & 20

OpenStudy (tkhunny):

Please give names to the two numbers.

OpenStudy (cenaida):

@him1618

OpenStudy (***[isuru]***):

Hi, Let's name the 2 numbers as A and B According to the info given, A + B = 24 -----------------(1) A - B = 14 ----------------(2) Now let's just solve this system , (1) + (2) A + B + ( A - B ) = 38 2A = 38 A = 19 by substituting A = 19 in (1) B = 24 -19 = 5 yep that's it the 2 numbers r 5 , 19 So, the choice is A Hope this will help ya out !!!

OpenStudy (cenaida):

Thank you so much @***[ISURU]***